To 130 16 MA=TABX FC Determine the Moment et a force about Point A. Express the result in cartesian vector form.

F-150 lb
Answer #1

A(4,5,0), B(1,0,14),c(-2,3,-2) AB=OB-OA=-3i – 5j +14k BC = OC - OB=-3i+3j –12k Fgc = 130/ -3i+37-12k F%c = 130 132 +32 +12) F